# Detecting Metadata-Related Bugs in Enterprise Applications

### v1

@@ 8.1. Rule Importance and Bug Criticality
The 15 rules we investigated are important, because we extracted them from the documentation of Spring and JUnit. All the real bugs we revealed using those rules are critical and severe, as they all trigger runtime instead of compilation errors. It means that without any tool support, developers have to wait till the testing phase, to reveal those rule violations. Additionally, developers spent lots of time revealing some of the bugs we found. As mentioned in Section 6.3, 15 of the bugs we detected were fixed 1-10 days after the bug-introducing commits; 17 of the bugs were fixed more than 10 days after the bug-introducing commits. The phenomena imply the difficulty of revealing those bugs, and the necessity of our approach.
@@ 9.1. Domain-Specific Languages (DSLs) Defined for Metadata-Usage Checking
Song and Tilevich created (1) MIL to express metadata invariants, and (2) the language implementation/engine to examine code-annotation relations (i.e., relations between Java code and annotations) and code-XML relations (i.e., relations between code and XML) (Song and Tilevich, 2012). We tried to run MIL engine in our experiments, albeit without success. RSL is different from MIL in three ways. First, it is an imperative instead of declarative language, so users have more flexibility in controlling how items are extracted, enumerated, filtered, and checked. Second, RSL has stronger expressiveness. It can express rules to examine XML-annotation-code constraints and XML-annotation constraints, but we did not find MIL capable of doing that. Third, RSL can have better performance, because we applied specialized optimizations in MeCheck to eliminate repetitive computation but MIL engine was unoptimized.
@@ 9.3. Configuration Debugging
Some tools were built to diagnose or fix software configuration errors (Attariyan and Flinn, 2008, 2010; Rabkin and Katz, 2011; Zhang and Ernst, 2013; Weiss et al., 2017; Oh et al., 2021). For instance, Attariyan et al. (Attariyan and Flinn, 2008) and Zhang et al. (Zhang and Ernst, 2013) separately created tools to record predicates that may be affected by configuration options, collect the execution profiles of a program’s correct and undesired runs, and compare the behavioral differences between the two types of runes to diagnose configuration errors. Rabkin and Katz created a static analysis-based tool to help users debug configuration errors in software (Rabkin and Katz, 2011). The tool tracks the flow of configuration labels through a program: labels are introduced via configuration reads, and propagate via assignment and library calls. MeCheck produces a table mapping each line in the program’s source code to the set of relevant configuration dependencies. This procedure is done by developers at release time. When a user encounters an error, they can use the error message to query this table.
Weiss et al. built an approach to generalize system configuration repairs for certain types of machines, from the shell commands developers entered to update one machine (Weiss et al., 2017). Oh et al. introduce a model checking framework for building Kconfig static analysis tools (Oh et al., 2021). The configuration specification language, Kconfig, is defined to prevent invalid configurations of the Linux kernel from being built. To detect bugs in Kconfig specifications, Oh et al. created a symbolic evaluator kclause to model Kconfig specifications, and a tool to find bugs in kclause models.
The configuration files examined by these tools are irrelevant to XML documents or Java annotations, so they cannot detect the bugs we focus on.
@@ 10. Conclusion
As with proper code implementation, correct metadata usage is always essential to software quality. Albeit the importance of high-quality metadata usage, widely used program analysis techniques (e.g., WALA (wal, 2024) and Soot (soo, 2024)) barely help with metadata debugging. Some existing tools can help developers identify certain metadata bugs, but the tool support is quite limited. In this paper, we developed RSL—a domain-specific language for describing metadata-usage checking rules, and built MeCheck—a tool to analyze programs based on RSL rules. Compared with existing tools, our approach is unique in three aspects. First, RSL enables library/framework developers to freely specify domain-specific constraints on the relations among code, annotations, and XML items; thus, it empowers MeCheck to enforce diverse rules in a wider scope and equips MeCheck with a strong power extensibility. Second, MeCheck extracts program data from diverse software artifacts (e.g., Java source files and XML files), and reasons about the relationship between the extracted data to detect bugs. It demonstrates a great way of integrating source code analysis with metadata analysis. Third, MeCheck applies optimizations to reduce redundant data processing, when trying to establish pattern matching between a given program and multiple bug patterns described by RSL rules. No prior work pays such a close attention to the time cost of metadata-related bug detection.
@@
11. Data Availability
Our programs and data are available at https://github.com/mahirkabir/detecting-metadata-bugs
@@ Acknowledgments
We thank all reviewers for their valuable feedback. This work was partially funded by NSF-2006278, NSF-1929701, NSF-186467, NSF-2007718, and CCI.

### v2

@@ 8.1. Rule Importance and Bug Criticality
The 15 rules we investigated are important, because we extracted them from the documentation of Spring and JUnit. All the real bugs we revealed using those rules are critical and severe, as they all trigger runtime instead of compilation errors. It means that without any tool support, developers have to wait till the testing phase, to reveal those rule violations. Additionally, developers spent lots of time revealing some of the bugs we found. As mentioned in Section 6.3, 15 of the bugs we detected were fixed 1-10 days after the bug-introducing commits; 17 of the bugs were fixed more than 10 days after the bug-introducing commits. These phenomena imply the difficulty of revealing those bugs, and the necessity of our approach.
@@ 9.1. Domain-Specific Languages (DSLs) Defined for Metadata-Usage Checking
Song and Tilevich created (1) MIL to express metadata invariants, and (2) the language implementation/engine to examine code-annotation relations (i.e., relations between Java code and annotations) and code-XML relations (i.e., relations between code and XML) (Song and Tilevich, 2012). We tried to run MIL engine in our experiments, but without success. RSL is different from MIL in three ways. First, it is an imperative instead of declarative language, so users have more flexibility in controlling how items are extracted, enumerated, filtered, and checked. Second, RSL has stronger expressiveness. It can express rules to examine XML-annotation-code constraints and XML-annotation constraints, but we did not find MIL capable of doing that. Third, RSL can have better performance, because we applied specialized optimizations in MeCheck to eliminate repetitive computation but MIL engine was unoptimized.
@@ 9.3. Configuration Debugging
Some tools were built to diagnose or fix software configuration errors (Attariyan and Flinn, 2008, 2010; Rabkin and Katz, 2011; Zhang and Ernst, 2013; Weiss et al., 2017; Oh et al., 2021). For instance, Attariyan et al. (Attariyan and Flinn, 2008) and Zhang et al. (Zhang and Ernst, 2013) separately created tools to record predicates that may be affected by configuration options, collect the execution profiles of a program’s correct and undesired runs, and compare the behavioral differences between the two types of runs to diagnose configuration errors. Rabkin and Katz created a static analysis-based tool to help users debug configuration errors in software (Rabkin and Katz, 2011). The tool tracks the flow of configuration labels through a program: labels get introduced via configuration reads, and propagated via assignment and library calls. Weiss et al. built an approach to generalize system configuration repairs for certain types of machines, from the shell commands developers entered to update one machine (Weiss et al., 2017). Oh et al. introduce a model checking framework for building Kconfig static analysis tools (Oh et al., 2021). The configuration specification language, Kconfig, is defined to prevent invalid configurations of the Linux kernel from being built. To detect bugs in Kconfig specifications, Oh et al. created a symbolic evaluator kclause to model Kconfig specifications, and a tool to find bugs in kclause models. The configuration files examined by these tools are irrelevant to XML documents or Java annotations, so they cannot detect the bugs we focus on.
@@ 10. Conclusion
Similar to proper code implementation, correct metadata usage is essential to software quality. Despite the importance of high-quality metadata usage, widely used program analysis techniques (e.g., WALA (wal, 2024) and Soot (soo, 2024)) provide little support for metadata debugging. Some existing tools can help developers identify certain metadata bugs, but the tool support is quite limited. In this paper, we developed RSL—a domain-specific language for describing metadata-usage checking rules, and built MeCheck—a tool to analyze programs based on RSL rules. Compared with existing tools, our approach is unique in three aspects. First, RSL enables library/framework developers to freely specify domain-specific constraints on the relations among code, annotations, and XML items; thus, it empowers MeCheck to enforce diverse rules in a wider scope and equips MeCheck with a strong power extensibility. Second, MeCheck extracts program data from diverse software artifacts (e.g., Java source files and XML files), and reasons about the relationship between the extracted data to detect bugs. It demonstrates a great way of integrating source code analysis with metadata analysis. Third, MeCheck applies optimizations to reduce redundant data processing, when trying to establish pattern matching between a given program and multiple bug patterns described by RSL rules. No prior work pays such a close attention to the time cost of metadata-related bug detection.
@@
Data Availability
The data and programs are available at https://doi.org/10.5281/zenodo.15205192.
@@ Acknowledgments
We thank all reviewers for their valuable feedback. This work was partially funded by NSF-2006278, NSF-1929701, NSF-1845446, NSF-2007718, and CCI.